### v1

@@ 3.2 Compressing Clean Models
Table 2: Compressed size for various Models with ZipNN method with Byte Grouping. Based on compressing 1GB from the middle of a model (for large models), and the entire model, excluding the first 10MB (headers etc…) for smaller models.
@@
Identifying incompressible data in advance has the benefit of speeding up compression time. Various methods have been discussed in prior work to achieve this, e.g. Harnik et al. (2013). We propose a simple method that builds on the fact that within each byte group compressibility seems to be consistent. We compress a chunk, and if it is not compressible we skip the compressibility of the following few chunks before trying to compress again (the specific number varies with the actual implementation). The rationale in retrying to compress is to identify a possible change in behavior between different layers or parts of the model.
3.3 Compressibility summary
Table 2 presents the compressibility of various models using ZipNNwith a breakdown of compressibility of the various byte groups. The first byte group consists of the exponent. We see all BF16 models achieve a space saving of about $\frac{1}{3}$ of the model size. Regular models with parameter types FP32 or FP16 show less impressive space savings whereas clean models of these type show much better space savings. In clean FP32 models byte grouping plays a key role as seen in the breakdown. We also see clean models in the FP16 family that are likely the result of transformation from BF16 models (which has a shorter fraction part than FP16).
@@ 3.3 Compressibility summary
4 Beyond Model compression
@@
4.1 Gradients and Optimizers
Gradients and Optimizers are derivatives of the training process. These artifacts are required in order to continue the training process and as such take up substantial network bandwidth (in the case of distributed training) or storage space (in the case of checkpoints). Gradients and Optimizers are often of equal in size to the models Anil et al. (2020) and therefore can also benefit from compression.
@@ 4.1 Gradients and Optimizers
4.2 Checkpoints and Delta Compression
When models have high similarity, one strategy to optimize storage and network transfer is to save a base model and for the rest of the models only store the differences from this base model Kandpal et al. (2023). We refer to compressing those differences as delta compression. To reconstruct a model, one only needs to apply the delta to the base model. A straightforward approach to delta compression is to compute the difference between the two models (e.g. using XOR or subtraction) and compress this delta using a lossless compressor. In our work, we used XOR for the delta as it is easily reversible and does not require any extra bits. A natural use case in which delta compression proves very useful is checkpointing. In checkpointing, we repeatedly store models that have limited change between them. Fine tuning often changes models in small quantities. Hence the delta between the results of consecutive training epochs has the potential to be more compressible than stand-alone models. We studied this while finetuning a Resnet 18 model (FP32). Figure 8(a) shows the amount of change between consecutive checkpoints as the process progresses and the loss function drops. We see that while all parameters in the model change in each epoch, when broken down to bytes, more and more bytes remain unchanged as the training converges. Figure 8(b) breaks this down according to byte groups. Here we see that the exponent byte has the least changes, whereas the least bits in the fraction have the most change. All byte groups show a growing number of steady bytes as the training converges. This suggests that byte grouping is helpful for delta compression as well, as our tests in Figure 8(c) corroborate.
@@ 4.2 Checkpoints and Delta Compression
5 Implementation and Evaluation
5.1 Implementation
We implemented our method and it is now an open source project ZipNN . It’s core is written in C (2000 lines of code) and the wrappers, scripts and tests are all written in Python (4000 lines of code). We use the Zstd v1.5.6 library as well as Zstd’s underlying Huffman implementation as the underlying compressors. For our tests in Python, we use zstandard 0.23.0, torch 2.4.0, and numpy 2.1.0.
Chunking. When designing our method, we aimed to create an implementation that could process small chunks independently, allowing each chunk to be handled in parallel. This architecture is ideal for GPUs, which contain many cores that, while less powerful than CPU cores, excel in concurrent tasks. Our implementation includes two levels of granularity: chunk level and byte-group level within each chunk. By default, each chunk is of size 256KB. In the BF16 format byte-groups are sized at 128KB whereas in FP32 byte-groups are of size 64KB. Parallelism can be conducted at the chunk level, and we also parallelize the byte-group compression/decompression processes. Auto decisions on compression method can be done at a byte-group granularity.
@@ 5.1 Implementation
Hugging Face Integration. For smooth integration with Hugging Face’s Transformers library Wolf et al. (2020), we implemented a mechanism that automatically decompresses downloaded models, reorders symbolic links in the local cache, updates metadata, and removes the compressed files. We also included an option for manual compression or decompression of models in the local cache.
5.2 Compression and Decompression speed
We ran our tests on an Apple M1 Max machine with 10 cores and 64GB of RAM running macOS Sonoma 14.3. The tests run in a single process and on a single core. Table 3 shows the speed benefits of our method vs. vanilla compressors on 3 representative models. Two regular models (BF16 and FP32) and one clean model (FP32). The exponent-extraction and byte grouping carry a performance penalty, that is amplified by the fact that the compressor works harder if it actually manages to find some repetitions, yet the use of Huffman only encoding makes up for this and we manage to improve both on speed and on compression ratio simultaneously.
The variation in speeds between the various models is also explained by their compressibility. While in the regular FP32 model $\frac{3}{4}$ of the model is non compressible and mostly skipped, in the BF16 this is only $\frac{1}{2}$ and in the clean model even less than that. Note that we tested LZ4 and Snappy on these models and while they are faster than all methods, they gain zero compression savings.
@@ 5.2 Compression and Decompression speed
Table 3: Comparing speeds of compression methods on two regular models. Zstd was used with default configuration. Our method, ZipNN, consists of Exponent-Extraction and Huffman only compression. EE-Zstd denotes Exponent-Extraction with Zstd compression on the exponent. Based on 10 runs on 1GB from the middle of the model. The maximum standard deviation was 2%.
5.3 End-2-End Evaluation
@@
The end-2-end timing behavior is dictated by the time to compress/decompress the model and the time to upload/download it respectively. We measured timing with lossless compression on the 3 representative models tested in the previous section. Figure 10 shows the timing of upload and download speeds of the three models. Each test was run 10 times for the cached reads and 5 times for the $1^{st}$ timers. The variance was almost entirely due to the network time and this standard deviation is depicted in the graph. The actual compression and decompression time had very little variance. For example, for the xlm-RoBERTa model the average time for the decompression was 3.92 seconds with a standard deviation of 0.017.
@@
Figure 10: Download and upload times of 3 models using our compression vs. the non-compressed version.
6 Related Work
6.1 Model Compression
@@
In the literature, ”model-compression” is a field of its own, aimed at creating smaller models that mimic the original model. Model-compression is hence the name for a set of tools that aim to accelerate models, usually at inference, by reducing their size Choudhary et al. (2020). Under such conditions, a method is allowed to reduce the accuracy, and is judged on its tradeoff between size and performance. This differs from lossless compression which is supposed to return the model to its original state after decompression.
@@
There are four main methods to reduce model size in that manner Choudhary et al. (2020). Pruning LeCun et al. (1989); Hanson & Pratt (1988); Zhu & Gupta (2017) (sometimes referred to as sparsification; Ma et al. (2021)) where parts of the model are removed, dedicated training or network architecture Oktay et al. (2019), distillation Gou et al. (2021) or otherwise training a smaller model from a better model Haroush et al. (2019), and quantization Gholami et al. (2021). There are also methods that combine several of those Polino et al. (2018), including the only work we have found to propose compression, which it applies after two other model-compression steps Han et al. (2015).
@@
Out of the model-compression techniques, quantization Gray (1984) is perhaps the most popular. Quantization is a method that bins weight values to a more coarse granularity. By design it trades-off accuracy for space and speed and is hence a lossy compression technique. Quantization is bound to increase the amount of entropy per byte in a model but is not trying to push it to the limit. As such, quantized models can potentially be further compressed using lossless compression. We see mixed results: with some off-the-shelf quantized models that have been quantized with GTPQ Frantar et al. (2022) and AWQ Lin et al. (2023) we find that they are still compressible, with a compressed size between 85-91%. On the other hand, we found that other quantized models using GGUF do not compress at all.
6.2 Other Related Work
@@
Similar to delta compression some works analyze dimensionality Aghajanyan et al. (2021); Gueta et al. (2023) or save the deltas for actions such as compositionally Ilharco et al. (2022) and merging multiple deltas Choshen et al. (2022); Wortsman et al. (2022); Matena & Raffel (2021). Few works also apply on such deltas methods like pruning Yadav et al. (2023), trained sparsity Zhang et al. (2023) or quantization Dettmers et al. (2023) or discuss deltas.
@@
Another line of work worth mentioning is computation graph optimization Sabne (2020); Wu (2023). Such works reduce the computation graph and perform optimization there. It is noteworthy to contrast it to our work. such work compresses the size of the computation graph, which speeds computation, but does not change the model weights, and it is hence orthogonal to our work. To validate that, we compressed pyTorch Wu (2023) models before and after compilation, finding compression works similarly well.
Last, related to the optimization process (see §4.1), a few recent works offer to reduce the information passed in gradient updates hence making them faster, further overcoming the decrease in information per example by seeing more examples per second Tyagi & Swany (2023); Zhao et al. (2024).
7 Conclusion

### v2

@@ III-B Compressing Clean Models
TABLE II: Compressed size for various Models using ZipNN with Byte Grouping. Based on compressing 1GB from the middle of a model (for large models), and the entire model, excluding the first 10MB (headers etc…) for smaller models.
@@
Identifying incompressible data in advance has the benefit of speeding up compression time. Various methods have been discussed in prior work to achieve this, e.g. [43]. We propose a simple method that builds on the fact that within each byte group compressibility seems to be consistent. We compress a chunk, and if it is not compressible we skip the compressibility of the following few chunks before trying to compress again (the specific number varies with the actual implementation). The rationale in retrying to compress is to identify a possible change in behavior between different layers or parts of the model.
III-C Compressibility summary
Table II presents the compressibility of various models using ZipNN with a breakdown of compressibility of the various byte groups. The first byte group consists of the exponent. We see all BF16 models achieve a space saving of about $\frac{1}{3}$ of the model size. Regular models with parameter types FP32 or FP16 show less impressive space savings whereas clean models of these type show much better space savings. In clean FP32 models byte grouping plays a key role as seen in the breakdown. We also see clean models in the FP16 family that are likely the result of transformation from BF16 models (which has a shorter fraction part than FP16).
@@ III-C Compressibility summary
IV Beyond Model compression
@@
IV-A Gradients and Optimizers
Gradients and Optimizers are derivatives of the training process. These artifacts are required in order to continue the training process and as such take up substantial network bandwidth (in the case of distributed training) or storage space (in the case of checkpoints). Gradients and Optimizers are often of equal in size to the models [44] and therefore can also benefit from compression.
@@ IV-A Gradients and Optimizers
IV-B Checkpoints and Delta Compression
When models have high similarity, one strategy to optimize storage and network transfer is to save a base model and for the rest of the models only store the differences from this base model [2]. We refer to compressing those differences as delta compression. To reconstruct a model, one only needs to apply the delta to the base model. A straightforward approach to delta compression is to compute the difference between the two models (e.g. using XOR or subtraction) and compress this delta using a lossless compressor. In our work, we used XOR for the delta as it is easily reversible and does not require any extra bits. A natural use case in which delta compression proves useful is checkpointing. In checkpointing, we repeatedly store models that have limited change between them. Fine tuning often changes models in small quantities. Hence the delta between the results of consecutive training epochs has the potential to be more compressible than stand-alone models. Note that the main benefit in this use case is storage space rather than offloading from the GPU. Offloading from GPU can, for the most part, be done offline, and we want to refrain from spending extra GPU cycles for compression at the GPU. Hence we focus on CPU based compression for this task. We studied this while finetuning a Resnet 18 model (FP32). Figure 8(a) shows the amount of change between consecutive checkpoints as the process progresses and the loss function drops. We see that while all parameters in the model change in each epoch, when broken down to bytes, more and more bytes remain unchanged as the training converges. Figure 8(b) breaks this down according to byte groups. Here we see that the exponent byte has the least changes, whereas the least bits in the fraction have the most change. All byte groups show a growing number of steady bytes as the training converges. This suggests that byte grouping is helpful for delta compression as well, as our tests in Figure 8(c) corroborate.
@@ IV-B Checkpoints and Delta Compression
V Implementation and Evaluation
V-A Implementation
We implemented ZipNN and it is now an open source project [45]. It’s core is written in C (2000 lines of code) and the wrappers, scripts and tests are all written in Python (4000 lines of code). We use the Zstd v1.5.6 library as well as Zstd’s underlying Huffman implementation as the underlying compressors. For our tests in Python, we use zstandard 0.23.0, torch 2.4.0, and numpy 2.1.0.
Chunking. When designing ZipNN, we aimed to create an implementation that could process small chunks independently, allowing each chunk to be handled in parallel. This architecture is ideal for GPUs, which contain many cores that, while less powerful than CPU cores, excel in concurrent tasks. Our implementation includes two levels of granularity: chunk level and byte-group level within each chunk. By default, each chunk is of size 256KB. In the BF16 format byte-groups are sized at 128KB whereas in FP32 byte-groups are of size 64KB. Parallelism can be conducted at the chunk level and the byte-group compression/decompression processes. Auto decisions on compression method are done at a byte-group granularity.
@@ V-A Implementation
Hugging Face Integration. For smooth integration with Hugging Face’s Transformers library [46], we implemented a mechanism that automatically decompresses downloaded models, reorders symbolic links in the local cache, updates metadata, and removes the compressed files. We also included an option for manual compression or decompression of models in the local cache.
V-B Compression and Decompression Speed - Single Thread
Table III shows the speed benefits of ZipNN vs. vanilla compressors on 3 representative models. Two regular models (BF16 and FP32) and one clean model (FP32). Zstd was used with default configuration. ZipNN consists of Exponent-Extraction and Huffman only compression while EE-Zstd denotes Exponent-Extraction with Zstd compression on the exponent. This was based on 10 runs over 1GB from the middle of the model and the maximum standard deviation observed was 2%. The tests were run on an Apple M1 Max machine with 10 cores and 64GB of RAM running macOS Sonoma 14.3 using a single thread and on a single core.
The exponent-extraction and byte grouping carry a performance penalty, yet the use of Huffman only encoding makes up for this and we manage to improve both on speed and on compression ratio simultaneously. The variation in speeds between the various models is also explained by their compressibility. While in the regular FP32 model $\frac{3}{4}$ of the model is non compressible and mostly skipped, in the BF16 this accounts for only $\frac{1}{2}$ of the model and in the clean model even less than that. Note that we tested LZ4 and Snappy on these models and while they are faster than all methods, they gain zero compression savings.
TABLE III: Comparing speeds of compression methods.
@@ V-B Compression and Decompression Speed - Single Thread
Figure 10: Decompression throughput (a) and compression throughput (b) as a function of the number of threads used for compressing 100MB, 1GB, and 10GB across three models. (c) focuses on a Llama 3.1 while also varying the number of workers. The block size here is the size per worker.
V-C Multi-threading
We also implemented a multi-threaded version and tested it on a high-performance OpenShift cluster, utilizing a pod equipped with Intel Xeon Platinum 8480+ processors. The CPU features two NUMA nodes with a total of 224 cores (56 cores per socket, 2 sockets per NUMA node) and 2TB of DRAM. Figures 10(a) and 10(b) show the results of a single worker as a function of number of threads. We see that the size of the model is critical in utilizing many cores and decompressing 10GB reaches much better throughput (more than 45GBs) than compressing 1GB, let alone 100MB. The number of threads to reach peak decompression speed varies accordingly. The compression throughput is much lower and peaks at around 16 threads.
We further evaluated ZipNN with multiple workers to avoid running across NUMA nodes. Each NUMA node contains 56 threads (corresponding to all threads from a single socket). This multi-worker approach scales better than using a single worker as seen in Figure 10(c), reaching up to 80GB/s for decompression throughput and up to 13GB/s for compression with 16 workers and doing so while each worker gets a block size of as low as 100MB.
V-D End-2-End Evaluation
V-D1 Model Hubs
@@
The end-2-end timing behavior is dictated by the time to decompress the model with 8 threads and the time to download it. We measured timing with our compression method on the 3 representative models tested in the previous section. Figure 11 shows the timing of download speeds and also tests the effect on upload times to the cloud. Each test was run 10 times for the cached reads and 5 times for the $1^{st}$ timers. The variance was almost entirely due to the network time and this standard deviation is depicted in the graph. The actual compression and decompression time had little variance. For example, for the xlm-RoBERTa model the average time for the decompression was 3.92 seconds with a standard deviation of 0.017.
@@
Figure 11: Download and upload times of 3 models using our compression (with 8 threads) vs. the non-compressed version.
V-D2 Inference Serving System
Another use case we evaluated is a high-end OpenShift cluster for inference serving. We measured model loading times in a cloned version of such a cluster using the same pod configuration described in section V-C. The models are loaded from a high-performance clustered file system (having been previously downloaded from Hugging Face). The file system is connected to the pods via a Persistent Volume Claim (PVC) and exhibits a throughput of 8GB/s per worker. The test was run using vLLM version 0.7.2 with 4 workers, loading a 16GB Granite-3.1-8b-instruct BF16 model [34] which had been compressed to $\frac{2}{3}$ of its original size. Using our multi-threaded implementation, loading the compressed model from PVC to CPU and then to GPU (after decompression) took approximately 3 seconds, on par with running using a non-compressed model. This demonstrates that we can store models in compressed format on the PVC to save storage space with negligible impact on pod initialization time.
VI Related Work
VI-A Model Compression
@@
In the literature, ”model-compression” is a field of its own, aimed at creating smaller models that mimic the original model. Model-compression is hence the name for a set of tools that aim to accelerate models, usually at inference, by reducing their size [47]. Under such conditions, a method is allowed to reduce the accuracy, and is judged on its tradeoff between size and performance. This differs from lossless compression which is supposed to return the model to its original state after decompression.
@@
There are four main methods to reduce model size in that manner [47]. Pruning [48, 49, 50] (sometimes referred to as sparsification; [8]) where parts of the model are removed, dedicated training or network architecture [51], distillation [7] or otherwise training a smaller model from a better model [52], and quantization [9]. There are also methods that combine several of those [53, 54].
@@
Out of the model-compression techniques, quantization [55] is perhaps the most popular. Quantization is a method that bins weight values to a more coarse granularity. By design it trades-off accuracy for space and speed and is hence a lossy compression technique. Quantization is bound to increase the amount of entropy per byte in a model but is not trying to push it to the limit. As such, quantized models can potentially be further compressed using lossless compression. We see mixed results: with some off-the-shelf quantized models that have been quantized with GTPQ [56] and AWQ [57] we find that they are still compressible, with a compressed size between 85-91%. On the other hand, we found that other quantized models using GGUF do not compress at all.
VI-B Floating Point Compression
There is a body of literature on compression tailored specifically for floating-point data type. Some compression methods [58, 59] propose separating the ”sign exponent” from the mantissa and compressing it independently as part of a more complex approach.
dietgpu [60] is an open source project that implements ANS compression in the GPU aiming to reduce incoming traffic to a GPU. One of their variants implements exponent only compression for floating point streams is similar to ZipNN on regular models. This falls short on clean models, checkpoint deltas, gradients and optimizers.
Other compression methods, such as ZFP [61], based on [62, 63], take a block-based approach by dividing data into small blocks. Within each block, floating-point values are converted to a block-floating-point format, sharing a single exponent while being transformed into signed integers. Compression is then applied to this representation, making it particularly effective for lossy compression with controllable error bounds. Our benchmarks using ZFP version 1.3.0 with FP32 show its compression ratio comparable to ZSTD but lower than ZipNN’s compression ratio.
VI-C Other Related Work
@@
Similar to delta compression, some works analyze dimensionality [64, 65] or save the deltas for actions such as compositionally [66] and merging multiple deltas [67, 68, 69]. Few works also apply on such deltas methods like pruning [70], trained sparsity [71] or quantization [72].
@@
Another line of work worth mentioning is computation graph optimization [73, 74]. Such works reduce the computation graph and perform optimization there. It is noteworthy to contrast it to our work. such work compresses the size of the computation graph, which speeds computation, but does not change the model weights, and it is hence orthogonal to our work. To validate that, we compressed pyTorch [74] models before and after compilation, finding compression works similarly well. Last, related to the optimization process (see §IV-A), a few recent works offer to reduce the information passed in gradient updates hence making them faster, further overcoming the decrease in information per example by seeing more examples per second [75, 76].
VII Conclusion